OpenStudy (anonymous):

Given the following perfect square trinomial, fill in the missing term. x2 − 16x + ____

OpenStudy (anonymous):

@xiphoidx

OpenStudy (cwrw238):

use the identity:- (x - a)^2 = x^2 -2a + a^2

OpenStudy (anonymous):

what's the answer? i don't really get this

OpenStudy (cwrw238):

compare the Right hand side to your trinomial

OpenStudy (anonymous):

x^2-2*x*8+8^2-8^2 =(x-8)^2-8^2 ans

OpenStudy (cwrw238):

if we compare term by term we get -2a = -16 so a^2 = ?

OpenStudy (anonymous):

is that the answer?

OpenStudy (cwrw238):

find the value of a from the two equations -2a = -16 so a = ? now fit this value of a into a^2 and you have your answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

a = -8?

OpenStudy (anonymous):

no...a=8 not -8

OpenStudy (anonymous):

oh yeah I always get that mixed up thanks

OpenStudy (cwrw238):